OpenStudy (anonymous):

in what form is the following linear equation written y-3=2/3(x-1)

OpenStudy (anonymous):

point slope formm..

OpenStudy (anonymous):

Thank you! ^.^

OpenStudy (anonymous):

compare it with this... (y-y1)=m(x-x1) here m=slope...

OpenStudy (anonymous):

you welcome dear..